OpenStudy (anonymous):

Solve the following equation by clearing fractions. Please show all work. 8/9-x/3 =2x-2

OpenStudy (anonymous):

\[\frac{8}{9}-\frac{x}{3}=2x-2\] \[9(\frac{8}{9}-\frac{x}{3})=9(2x-2)\] \[8-3x=18x-18\] \[8+18=18x+3x\] \[26=21x\] \[x=\frac{26}{21}\]